OpenStudy (anonymous):

Find the point on the line -3 x + 3 y + 5 =0 which is closest to the point ( -5, 2 ).

OpenStudy (jdoe0001):

hmm have you covered slopes yet?

OpenStudy (anonymous):

I know that f '(2) = (19e^6)/98. Then after I know I have to switch the slopes somehow to evaluate where they intersect but I'm getting wrong values. I get lost after this step.

OpenStudy (freckles):

you need to minimize the distance

OpenStudy (freckles):

that is differentiate the distance

OpenStudy (freckles):

solve -3x+3y=-5 for y

OpenStudy (freckles):

a point on -3x+3y=-5 is (x,y) we are solve -3x+3y=-5 for y because we want to put our point (x,y) just in terms of one variable

OpenStudy (anonymous):

y=-5/3 - x. I got that, then Im supposed to flip the slope y = -1/x..?

OpenStudy (freckles):

is this a calculus question or not? I thought it was when I seen f' above but now that I think about it your f' is really weird

OpenStudy (freckles):

aslo y=x-5/3 isn't it?

OpenStudy (freckles):

i must go for now